(1 point) Find the angle a between the vectors 4 -2 -2 3 and -3 3 ณ = 2.8
Answer #1

-2 J Sel) het ü y: 3 -3 -2 3 G = (-2)(4) + (3)(-3)+(C)(3) = -23 Q || |-)+(3) + ly TING | = (4) + (-3Y+ (3) = J39 - E 反 The an